OpenStudy (moongazer):

Find the equation of a line passing trough point (3,-2) and having a slope of -1.

OpenStudy (anonymous):

y-(-2)=-1(x-3) y +2 = -x+3 y = -x +3-2 y = -x+1 If I deserve medal then give me one

OpenStudy (moongazer):

so why did you do that? please explain ill give you a medal if you explained it well

OpenStudy (anonymous):

\[y-y_{1}=m(x-x_{1})\]\[y_{1}=-2\]\[x_{1}=3\] slope = -1 given and solve

OpenStudy (anonymous):

m=slope=-1